Reports 1-1 of 1 Clear search Modify search
AdV-DAQ (Calibration)
lagabbe, spinicelli - 17:14 Wednesday 22 September 2021 (53227) Print this report
Laser shutted on WE PCal, in front of photodiode Tx_PD1

The 22/09, at 9h06 ut, Piernicola shutted the laser beam manually, in front of the photodiode PCAL_WE_Tx_PD1. In order to measure the frequency response of the photodiode.

Images attached to this report
Comments to this report:
lagabbe - 11:14 Friday 24 September 2021 (53258) Print this report

Same the 23/09 at 7h05 UT,

Images attached to this comment
Search Help